# MedHal: An Evaluation Dataset for Medical Hallucination Detection

### v1

@@ 3.2 Question Answering Dataset Transformation
3.3 Information Extraction Dataset Transformation
Information extraction (IE) datasets comprise, for each source document (e.g., clinical notes, clinical trials), a set of extracted text sequences related to specific concepts within that document. These extracted sequences, or ”extractions,” represent information pertaining to a particular attribute of the document. For example, from a clinical note, an extraction might represent the patient’s reason for the visit. To generate factual samples, the source document is used as the context, and the corresponding extraction is treated as the statement. Non-factual samples are generated by randomly interchanging extractions of the same value type between different documents (e.g., medications are swapped between patient records). When extractions are structured as key-value pairs, a large language model (LLM) is employed to transform these pairs into declarative statements. For instance, the key-value pair ”visit motivation: Lower back pain” is converted into the statement ”The patient’s visit motivation is lower back pain”. This transformation is performed using the prompt template shown in the Appendix 4 within a one-shot prompting framework. The explanation for a non-factual statement is provided by the factual extraction associated with the same document and key concept. An example of how a sample from an information extraction dataset is transformed is shown in Table 3. In cases where extractions are related to attributes with limited value diversity, such as sex, we ensure distinct values by explicitly switching them if random swapping results in identical values. The source documents are used without any preprocessing.
@@ 3.3 Information Extraction Dataset Transformation
3.4 Natural Language Inference Dataset Transformation
3.5 Summarization Dataset Transformation
For summarization datasets, factual samples are constructed using genuine summaries or sentences extracted from them. Hallucinated samples are generated by extracting a single sentence from the original summary and employing an LLM to modify it, introducing self-contradictory information. The modified, hallucinated sentence is then reintegrated into the summary at its original position. The source text intended for summarization serves as the context, and the summary itself is treated as the statement. The explanation for a non-factual statement is simply the original sentence used to generate the hallucinated version. Sentences for hallucination generation are selected randomly, with a minimum length of 100 characters to ensure sufficient contextual information for the LLM. An example of how a sample is generated is shown in Table 4. The prompt used to generate self-contradictory sentences is shown in Appendix 5.The goal of this process is to evaluate if models can effectively detect small mistakes in statements consisting of long sequences of text.
@@ 3.5 Summarization Dataset Transformation
3.6 Dataset Description
Dataset Task S Content Type Source # Samples # Gen
MedMCQA QA Medical Content Pal et al. (2022) 183,000 70,730
MedNLI NLI Clinical Notes Herlihy & Rudinger (2021) 11,232 7,488
ACM IE Clinical Notes Bonnet & Boulenger (2024) 22,000 73,040
MedQA QA Medical Content Jin et al. (2020) 12,723 18,906
PubMedSum Sum Clinical Trials Gupta et al. (2021) 33,772 178,657
Table 5: Description of datasets used to generate the MedHal benchmark (S is for Synthetic)
The discrepancy between the original sample count and the generated sample count is attributable to several factors. Firstly, for certain tasks like information extraction, original samples can yield multiple generated samples. Secondly, to manage computational resource limitations, the total number of generated samples was reduced. Finally, samples resulting in token sequences longer than 8192 were excluded to mitigate potential adverse effects on model training. Preprocessing was selectively applied to specific datasets. For the ACM dataset, only valid JSON-structured extractions were retained. For QA datasets, explanations containing phrases such as ”Answer is c)” that referred to the original dataset options were removed using regular expressions, as we are using these samples outside the QA setting. For instance, the explanation ”Answer is c), Metformin is most commonly used as a first-line treatement, because…)” would be transformed to ”Metformin is most commonly used as a first-line treatement, because…)”.
The MedHal benchmark is partitioned into training, validation, and test sets according to an 80/10/10 ratio, ensuring a consistent evaluation framework. This results in 313,920 samples for the training set, 17458 samples in the validation set and 17443 samples in the test set.
4 Experiments
4.1 Models
For the generation of the MedHal dataset, we utilize the Llama-3-70B model (Grattafiori et al. (2024)), a state-of-the-art language model known for its strong performance across various language generation tasks. For all prompting strategies employed in the sample generation process, a one-shot learning approach was adopted to provide the model with examples of the desired output format prior to generation. Importantly, to ensure a fair evaluation of our proposed benchmark, the dataset generation was performed exclusively using the training splits of the individual source datasets. This approach allows for the independent evaluation of our trained models on the original test sets of the constituent datasets.
4.2 Evaluator
To demonstrate the utility of the MedHal dataset, we fine-tuned a dedicated model for the detection of medical hallucinated content. This evaluator model is initialized with the Llama-1B model (Grattafiori et al. (2024)) and subsequently fine-tuned on the training set of the MedHal dataset. The detailed training configuration is provided in Appendix 7. The selection of a 1 billion parameter model as the base for our evaluator is motivated by the desire to provide a resource that is accessible to a wider range of researchers and domain experts, including those with limited computational resources. Existing evaluator models and general-purpose language models often comprise several billion parameters, necessitating significant infrastructure for their deployment. This can restrict the practical application of LLM-based evaluation metrics in certain research contexts. Furthermore, the use of a smaller model contributes to lower infrastructure costs associated with training and deployment. Future work will explore the potential for enhancing the evaluation metrics through the utilization of larger models or the optimization of the fine-tuning process.
4.3 Results
We also fine-tuned a general-purpose Llama-3 1B parameter model (MedHal-Llama-1B) on the dataset. This serves as a baseline to evaluate whether specialized medical LLMs, which have undergone extensive medical domain fine-tuning, already demonstrate strong performance in detecting hallucinated content and can effectively function as judges in medical hallucination detection tasks. The prompt used for all non fine-tuned models is shown in Appendix 6. We report two types of metrics: factuality metrics and explanation metrics. Factuality metrics aim to measure if the model correctly detected hallucinated content. These include the precision, recall, and F1-score. Explanation metrics aim to evaluate if explanations generated by models when statements are not factual are valid. They evaluate if models, when detecting hallucinated content, correctly identify which part of the statement is hallucinated. In order to compute these metrics, we only take into account samples where both the prediction and the ground truth indicate a non-factual statements since in other cases, a true explanation might not exist or not be generated. Explanation metrics include the ROUGE-1 (R1) and ROUGE-2 (R2) (Lin (2004)) scores and the BLEU (Papineni et al. (2002)) score. Results are shown in Table 6.
FT Type Models Factuality Explanation
P R F1 BLEU R1 R2
General Llama-1B-Instruct 0.53 0.32 0.40 0.01 0.13 0.03
Llama-8B-Instruct 0.59 0.62 0.60 0.05 0.24 0.14
Medical BioMistral-7B 0.56 0.43 0.49 0.03 0.22 0.11
MedLlama-8B 0.52 0.59 0.55 0.03 0.21 0.08
OpenBioLLM-8B 0.52 0.77 0.62 0.04 0.21 0.10
General MedHal-Llama (Ours) 0.75 0.77 0.76 0.45 0.70 0.59
Table 6: Results of models on test set of MedHal (FT: Fine-Tuned, P: Precision, R: Recall, F1: F1-Score, R1: ROUGE-1, R2: ROUGE-2)
As models do not always follow the output format given in the prompt in Appendix 6, we only compute the factuality metrics on samples that matched the output format. We also compute the F1-score of each model, based on the origin of each sample to see if models tend to excel in certain tasks compared to other tasks. These results are shown in Figure 2.
Figure 2: Accuracy of models based on each dataset source
5 Discussion
5.1 Overall Model Performance
Our evaluation reveals several key insights regarding the performance of various models in detecting and explaining medical hallucinations. Notably, general-purpose models demonstrated a surprising aptitude for identifying factual inaccuracies, with Llama-8B-Instruct achieving the highest precision (0.59) in models that were not fine-tuned. However, part of this performance might also be due to the memorization of samples available in our dataset as it is built from datasets training sets. We base this hypothesis on the observation that this model significantly outperformed others by a wide margin on MedNLI, the only dataset that had not been transformed by another LLM. In contrast, specialized medical models, despite extensive fine-tuning on medical texts, exhibited lower performance in factuality metrics. Only OpenBioLLM demonstrated higher overall performance with an F1-score of 0.62. However, it exhibited a higher recall than precision, indicating a tendency towards overconfidence in identifying factual statements. Interestingly, general-purpose models also outperformed medical models in generating explanations for identified hallucinations, as evidenced by higher BLEU and ROUGE scores. After manual evaluation, this is also mainly due to medical models not generally following instructions to explain their choices.
5.2 Task-Specific Performance Variations
Performance varied across different tasks, highlighting the challenges associated with specific data sources. Notably, information extraction proved challenging for all models, including our fine-tuned MedHal-Llama-1B, which performed only slightly better than random chance. We hypothesize that this might be attributable to the synthetic nature of the information extraction dataset, which might introduce some inaccuracies in our labels. Furthermore, we observed no significant performance disparity between tasks that require contextual analysis (information extraction, summarization, NLI) and tasks that operate on isolated statements (question-answering). This suggests that the presence or absence of contextual information does not substantially impact the models’ ability to detect and explain hallucinations.
5.3 Fine-Tuned Evaluation Model
Fine-tuning a 1 billion parameter Llama model (MedHal-Llama) on our dataset yielded impressive results, surpassing the performance of models seven to eight times its size, including those with extensive medical fine-tuning. This finding might indicate that specialized medical fine-tuning does not guarantee a higher performance in hallucination detection, at least as measured by our dataset. MedHal-Llama especially demonstrated high performance in detecting hallucinations within the summarization task in both factuality and explanation metrics. This is likely due to the overrepresentation of summarization samples in our dataset, leading the model to correctly identifying which sentences were falsified. Future research should however focus on balancing the dataset across all tasks to ensure a more equitable evaluation.
6 Limitations and Future Work
It is important to acknowledge some limitations of our methodology, the main one being that we rely on generated content. The validity of the generated hallucinated content was primarily assessed through visual inspection of a sample of the generated data and relied on the performance of a 70 billion model at not hallucinating over small sequences of texts. Indeed, all generations only involved generating a small number of tokens for the model. However, ideally, a comprehensive evaluation involving medical professionals would be conducted to validate the accuracy and clinical relevance of the generated dataset. Additionally, while we only fine-tune a general purpose model, future work should also evaluate the performance of fine-tuning medical models to this task to see how they compare to our current fine-tuned model that is based on a general purpose model.
7 Conclusion
In this study, we introduced MedHal, the first large-scale dataset designed to specifically address the challenge of hallucination detection across diverse clinical tasks, including summarization, natural language inference, information extraction, and question answering. We evaluated a range of medical and general-purpose models on this dataset, providing a comprehensive benchmark for assessing their performance in this critical domain. Furthermore, we fine-tuned an evaluator model on MedHal, which we make available to the community as a baseline and tool for future research. While MedHal represents a significant step forward in addressing the limitations of existing hallucination detection methods in the medical field, we acknowledge the need for further refinement. Specifically, future work should prioritize validating the accuracy and clinical relevance of labels and explanations through expert reviews. Additionally, it should also focus on balancing the dataset across all included tasks to ensure a more equitable evaluation. By addressing these limitations, we can enhance the reliability and utility of MedHal.
8 Acknowledgements
We gratefully acknowledge the financial support provided by the FRQS (Fonds de Recherche du Québec – Santé) Chair in Artificial Intelligence and Digital Health. This research also benefited from the computational resources and GPU infrastructure provided by the Digital Research Alliance of Canada (DRAC) and Calcul Québec. We extend our sincere thanks to Prof. John Kildea for his valuable input during initial discussions on the challenges of working in clinical settings.
Appendix A Appendix
@@
Figure 4: Prompt template used for the IE task (for more information, see Bonnet & Boulenger (2024))
@@
### Task Description - You will evaluate whether a medical statement is factually accurate. - The statement may reference a provided context. - Respond with ”YES” if the statement is factually correct or ”NO” if it contains inaccuracies. - In order to answer YES, everything in the statement must be supported by the context. - In order to answer NO, there must be at least one piece of information in the statement that is not supported by the context. - You must also provide an explanation of why you think the statement is factual or not. If it is factual, put ”The statement is factual” as your explanation. - Your answer should follow the following format : Factual: [YES/NO] Explanation: [Your explanation] ### Context [context] ### Statement [statement]

### v2

@@ Appendix A Data Transformation Examples
Appendix B Prompt templates
@@
Figure 3: Prompt template used for the IE task (for more information, see Bonnet and Boulenger (2024))
@@
### Task Description - You will evaluate whether a medical statement is factually accurate. - The statement may reference a provided context. - Respond with "YES" if the statement is factually correct or "NO" if it contains inaccuracies. - In order to answer YES, everything in the statement must be supported by the context. - In order to answer NO, there must be at least one piece of information in the statement that is not supported by the context. - You must also provide an explanation of why you think the statement is factual or not. If it is factual, put "The statement is factual" as your explanation. - Your answer should follow the following format : Factual: [YES/NO] Explanation: [Your explanation] ### Context [context] ### Statement [statement]
@@
### Task Description - You will evaluate whether a medical statement is factually accurate. - The statement may reference a provided context. - Respond with "YES" if the statement is factually correct or "NO" if it contains inaccuracies. - In order to answer YES, everything in the statement must be supported by the context. - In order to answer NO, there must be at least one piece of information in the statement that is not supported by the context. - You must also provide an explanation of why you think the statement is factual or not. If it is factual, put "The statement is factual" as your explanation. ### Context [context] ### Statement [statement]
Figure 6: Prompt template used for training models
Appendix C Training configuration
@@
Appendix D Annotation Guidelines
Evaluation Criteria
What Constitutes a “Factual” Statement?
A statement is considered factual if and only if:
  1. 1.
    Every piece of information mentioned in the statement can be directly supported by the provided context, OR
  2. 2.
    When no context is provided, the statement represents well-established, accurate medical knowledge.
  3. 3.
    There are no contradictions between the statement and the context.
  4. 4.
    The statement does not include unsupported claims or extrapolations beyond what the context states.
Evaluation Process
  1. 1.
    Review the Context
    Carefully read the provided context (e.g., research abstract, clinical information, etc.)
  2. 2.
    Analyze the Statement
    Break down the statement into individual claims and check each claim against the context.
  3. 3.
    Cross-Reference with Medical Knowledge
    For statements without context, evaluate them against established medical facts.
  4. 4.
    Make Your Assessment
    Fill in the evaluation columns:
    • Valid Column:
      • YES: You agree with the current label (TRUE for factual / FALSE for non-factual)
      • NO: You disagree with the current label
    • Comment Column: Provide your reasoning, including:
      • Specific discrepancies found (if any)
      • Medical knowledge that supports or contradicts the statement
      • Concerns about accuracy or completeness
Example Evaluation
Context:
“A randomized controlled trial of 200 patients found that Drug A reduced hospital readmissions by 30% compared to placebo (p<0.05p<0.05).”
Statement:
“Drug A significantly reduces hospital readmissions by 30% in all patients.”
Current Label: TRUE
Your Evaluation:
  • Valid: NO
  • Comment: “While the 30% reduction is accurate, the statement overgeneralizes by claiming effectiveness ‘in all patients’ when the study only included 200 patients with specific characteristics. The word ‘significantly’ is supported by p<0.05p<0.05.”
Appendix E Error Analysis
To uncover the underlying reasons why models fail to identify factual and non-factual statements, we perform an error analysis by categorizing each incorrect factuality prediction by its original task.
Figure 8: Proportion of Model Errors by Task Source
The results shown in Figure 8 reveal a consistent pattern: the vast majority of errors for most models originate from samples derived from the Question Answering task. This trend likely originates from the nature of these samples, which often assess general medical knowledge rather than requiring contextual analysis of a provided document. This suggests that while the models may possess the necessary reasoning capabilities to detect inconsistencies within a given text, they frequently lack the specific, ingrained medical knowledge required to validate standalone factual assertions. Suprisingly, this knowledge-versus-reasoning gap is mainly illustrated by OpenBioLLM-8B, a medical model that performs well on contextual tasks but whose errors are almost exclusively from the QA task. This profile suggests that its failures are not due to flawed logic but to a deficit in its medical knowledge. In contrast, a smaller model like Llama-3.2-1B displays a more uniform error distribution across all tasks, indicating more fundamental limitations in both its reasoning and knowledge faculties. Ultimately, this analysis indicates that a primary obstacle for many current small models is not a failure of reasoning itself, but insufficient medical knowledge, a weakness that is most exposed when contextual clues are absent and pure factual recall is demanded.
Additionally, we also investigate whether medical hallucination detection is mainly linked to context length. To do so, we plot the distribution of context lengths across samples that were not correctly labelled by models.
Figure 9: Distribution of context lengths across samples that were not correctly labelled by models in the test set
Figure 9 illustrates the frequency distribution of context lengths for samples where models did not correctly label the sample, relative to the overall distribution of context lengths in the test set. Several key trends emerge from this visualization. Firstly, OpenBioLLM seems to make mistakes on samples with shorter context lengths. This finding aligns with observations from Figure 8, as the Question Answering task typically involves shorter contexts compared to others. Second, several models (BioMistral-7B, Llama-3-8B, MedLlama-8B, Prometheus-2-8x7B, HallOumi-8B) have a higher peek density at the first mode than the second mode contrary to the test set distribution. This trend is particularly surprising, as models generally tend to struggle more with longer contexts. These results indicate that the task is a better indicator of the errors of a model than the actual context length. Only Llama-3.2-1B seems to have a distribution similar to the test set’s, indicating that context length does not have a big impact on performance. This might be due to current models supporting a larger context length. Indeed, our samples might not test to the limit the context retrieval capabilities of models.